OpenStudy (anonymous):

The amount of heat needed to raise 250 g of a substance by 15 K is 2,930 J. What is the specific heat of the substance (J/(kg*K))? A. 0.78 B. 11,000 C. 780 D. 1.1 × 10^7 **not 100% sure, but is this choice C? 780?

OpenStudy (anonymous):

 the equation... in case i didn't write it clearly :P

OpenStudy (michele_laino):

here we have to use this formula: \[\Large Q = cm\Delta T\] Q is the heat , c is the specific heat, and T is the absolute temperature

OpenStudy (anonymous):

okay! so what would be plugged in? :/

OpenStudy (michele_laino):

so, we can write: \[\Large c = \frac{Q}{{m\Delta T}} = ...\frac{{Joules}}{{Kg \times Kelvin}}\]

OpenStudy (anonymous):

okay! so 2930/(kg*15)? is that right? and solve for kg?

OpenStudy (michele_laino):

no, we have to compute this: \[\Large c = \frac{Q}{{m\Delta T}} = \frac{{2930}}{{0.250 \times 15}} = ...\frac{{Joules}}{{Kg \times Kelvin}}\]

OpenStudy (anonymous):

ohh okay!! so we get 781.333333 so about 780?

OpenStudy (michele_laino):

that's right!!

OpenStudy (anonymous):

yay!! thank you:D

OpenStudy (michele_laino):

:D